\begin{thebibliography}{10}
\providecommand{\url}[1]{\texttt{#1}}
\providecommand{\urlprefix}{URL }
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi:\discretionary{}{}{}#1}\else
  \providecommand{\doi}{doi:\discretionary{}{}{}\begingroup
  \urlstyle{rm}\Url}\fi
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem{EDELSTEIN1990233}
V.~Edelstein,
\newblock \emph{\textit{Spin polarization of conduction electrons induced by
  electric current in two-dimensional asymmetric electron systems}},
\newblock Solid State Commun. \textbf{73}, 233 (1990),
\newblock \doi{10.1016/0038-1098(90)90963-C}.

\bibitem{RevModPhys.91.035004}
A.~Manchon, J.~\ifmmode~\check{Z}\else \v{Z}\fi{}elezn\'y, I.~M. Miron,
  T.~Jungwirth, J.~Sinova, A.~Thiaville, K.~Garello and P.~Gambardella,
\newblock \emph{\textit{Current-induced spin-orbit torques in ferromagnetic and
  antiferromagnetic systems}},
\newblock Rev. Mod. Phys. \textbf{91}, 035004 (2019),
\newblock \doi{10.1103/RevModPhys.91.035004}.

\bibitem{9427163}
Q.~Shao, P.~Li, L.~Liu, H.~Yang, S.~Fukami, A.~Razavi, H.~Wu, K.~Wang,
  F.~Freimuth, Y.~Mokrousov, M.~D. Stiles, S.~Emori \emph{et~al.},
\newblock \emph{\textit{Roadmap of Spin–Orbit Torques}},
\newblock IEEE Trans. Magn. \textbf{57}, 1 (2021),
\newblock \doi{10.1109/TMAG.2021.3078583}.

\bibitem{PhysRevB.94.121114}
J.-P. Hanke, F.~Freimuth, A.~K. Nandy, H.~Zhang, S.~Bl\"ugel and Y.~Mokrousov,
\newblock \emph{Role of berry phase theory for describing orbital magnetism:
  From magnetic heterostructures to topological orbital ferromagnets},
\newblock Phys. Rev. B \textbf{94}, 121114 (2016),
\newblock \doi{10.1103/PhysRevB.94.121114}.

\bibitem{PhysRevResearch.3.013275}
A.~Johansson, B.~G\"obel, J.~Henk, M.~Bibes and I.~Mertig,
\newblock \emph{\textit{Spin and orbital Edelstein effects in a two-dimensional
  electron gas: Theory and application to ${\mathrm{SrTiO}}_{3}$ interfaces}},
\newblock Phys. Rev. Res. \textbf{3}, 013275 (2021),
\newblock \doi{10.1103/PhysRevResearch.3.013275}.

\bibitem{atencia2024orbital}
R.~B. Atencia, A.~Agarwal and D.~Culcer,
\newblock \emph{\textit{Orbital angular momentum of Bloch electrons:
  equilibrium formulation, magneto-electric phenomena, and the orbital Hall
  effect}},
\newblock Adv. Phys.: X \textbf{9}, 2371972 (2024),
\newblock \doi{10.1080/23746149.2024.2371972}.

\bibitem{doi:10.1021/acs.nanolett.7b04300}
T.~Yoda, T.~Yokoyama and S.~Murakami,
\newblock \emph{\textit{Orbital Edelstein Effect as a Condensed-Matter Analog
  of Solenoids}},
\newblock Nano Lett. \textbf{18}, 916 (2018),
\newblock \doi{10.1021/acs.nanolett.7b04300}.

\bibitem{salemi2019orbitally}
L.~Salemi, M.~Berritta, A.~K. Nandy and P.~M. Oppeneer,
\newblock \emph{\textit{Orbitally dominated Rashba-Edelstein effect in
  noncentrosymmetric antiferromagnets}},
\newblock Nature communications \textbf{10}, 5381 (2019),
\newblock \doi{10.1038/s41467-019-13367-z}.

\bibitem{PhysRevB.107.115402}
T.~P. Cysne, F.~S.~M. Guimar\~aes, L.~M. Canonico, M.~Costa, T.~G. Rappoport
  and R.~B. Muniz,
\newblock \emph{\textit{Orbital magnetoelectric effect in nanoribbons of
  transition metal dichalcogenides}},
\newblock Phys. Rev. B \textbf{107}, 115402 (2023),
\newblock \doi{10.1103/PhysRevB.107.115402}.

\bibitem{go2017toward}
D.~Go, J.-P. Hanke, P.~M. Buhl, F.~Freimuth, G.~Bihlmayer, H.-W. Lee,
  Y.~Mokrousov and S.~Bl{\"u}gel,
\newblock \emph{\textit{Toward surface orbitronics: giant orbital magnetism
  from the orbital Rashba effect at the surface of sp-metals}},
\newblock Sci. Rep. \textbf{7}, 46742 (2017),
\newblock \doi{10.1038/srep46742}.

\bibitem{PhysRevResearch.5.043294}
S.~Leiva-Montecinos, J.~Henk, I.~Mertig and A.~Johansson,
\newblock \emph{\textit{Spin and orbital Edelstein effect in a bilayer system
  with Rashba interaction}},
\newblock Phys. Rev. Res. \textbf{5}, 043294 (2023),
\newblock \doi{10.1103/PhysRevResearch.5.043294}.

\bibitem{PhysRevB.100.075136}
G.~Massarelli, B.~Wu and A.~Paramekanti,
\newblock \emph{\textit{Orbital Edelstein effect from density-wave order}},
\newblock Phys. Rev. B \textbf{100}, 075136 (2019),
\newblock \doi{10.1103/PhysRevB.100.075136}.

\bibitem{ishitobi2019magnetoelectric}
T.~Ishitobi and K.~Hattori,
\newblock \emph{\textit{Magnetoelectric Effects and Charge-Imbalanced
  Solenoids: Antiferro Quadrupole Orders in a Diamond Structure}},
\newblock J. Phys. Soc. Jpn. \textbf{88}, 063708 (2019),
\newblock \doi{10.7566/JPSJ.88.063708}.

\bibitem{levitov1985magnetoelectric}
L.~Levitov, Y.~V. Nazarov and G.~Eliashberg,
\newblock \emph{\textit{Magnetoelectric effects in conductors with mirror
  isomer symmetry}},
\newblock Sov. Phys. JETP \textbf{61}, 133 (1985).

\bibitem{PhysRevB.96.035120}
J.~Rou, C.~\ifmmode~\mbox{\c{S}}\else \c{S}\fi{}ahin, J.~Ma and D.~A. Pesin,
\newblock \emph{\textit{Kinetic orbital moments and nonlocal transport in
  disordered metals with nontrivial band geometry}},
\newblock Phys. Rev. B \textbf{96}, 035120 (2017),
\newblock \doi{10.1103/PhysRevB.96.035120}.

\bibitem{ishizaka2011giant}
K.~Ishizaka, M.~Bahramy, H.~Murakawa, M.~Sakano, T.~Shimojima, T.~Sonobe,
  K.~Koizumi, S.~Shin, H.~Miyahara, A.~Kimura \emph{et~al.},
\newblock \emph{\textit{Giant Rashba-type spin splitting in bulk BiTeI}},
\newblock Nat. Mater. \textbf{10}, 521 (2011),
\newblock \doi{10.1038/nmat3051}.

\bibitem{PhysRevB.90.195210}
L.~Wu, J.~Yang, S.~Wang, P.~Wei, J.~Yang, W.~Zhang and L.~Chen,
\newblock \emph{\textit{Two-dimensional thermoelectrics with Rashba spin-split
  bands in bulk BiTeI}},
\newblock Phys. Rev. B \textbf{90}, 195210 (2014),
\newblock \doi{10.1103/PhysRevB.90.195210}.

\bibitem{PhysRevB.55.16293}
E.~A. de~Andrada~e Silva, G.~C. La~Rocca and F.~Bassani,
\newblock \emph{\textit{Spin-orbit splitting of electronic states in
  semiconductor asymmetric quantum wells}},
\newblock Phys. Rev. B \textbf{55}, 16293 (1997),
\newblock \doi{10.1103/PhysRevB.55.16293}.

\bibitem{PhysRevLett.84.6074}
D.~Grundler,
\newblock \emph{\textit{Large Rashba Splitting in InAs Quantum Wells due to
  Electron Wave Function Penetration into the Barrier Layers}},
\newblock Phys. Rev. Lett. \textbf{84}, 6074 (2000),
\newblock \doi{10.1103/PhysRevLett.84.6074}.

\bibitem{Rashba}
E.~I. Rashba and V.~I. Sheka,
\newblock \emph{\textit{Symmetry of Energy Bands in Crystals of Wurtzite Type
  II. Symmetry of Bands with Spin-Orbit Interaction Included}},
\newblock Fiz. Tverd. Tela: Collected Papers \textbf{2}, 62 (1959),
\newblock \eprint{http://nrs.harvard.edu/urn-3:HUL.InstRepos:29426010}.

\bibitem{PhysRevLett.5.371}
R.~C. Casella,
\newblock \emph{\textit{Toroidal Energy Surfaces in Crystals with Wurtzite
  Symmetry}},
\newblock Phys. Rev. Lett. \textbf{5}, 371 (1960),
\newblock \doi{10.1103/PhysRevLett.5.371}.

\bibitem{PhysRevB.95.195164}
V.~Brosco and C.~Grimaldi,
\newblock \emph{\textit{Anisotropy of transport in bulk Rashba metals}},
\newblock Phys. Rev. B \textbf{95}, 195164 (2017),
\newblock \doi{10.1103/PhysRevB.95.195164}.

\bibitem{kubo1957statistical}
R.~Kubo,
\newblock \emph{\textit{Statistical-mechanical theory of irreversible
  processes. I. General theory and simple applications to magnetic and
  conduction problems}},
\newblock J. Phys. Soc. Jpn \textbf{12}(6), 570 (1957),
\newblock \doi{10.1143/JPSJ.12.570}.

\bibitem{PhysRevB.75.045315}
N.~A. Sinitsyn, A.~H. MacDonald, T.~Jungwirth, V.~K. Dugaev and J.~Sinova,
\newblock \emph{\textit{Anomalous Hall effect in a two-dimensional Dirac band:
  The link between the Kubo-Streda formula and the semiclassical Boltzmann
  equation approach}},
\newblock Phys. Rev. B \textbf{75}, 045315 (2007),
\newblock \doi{10.1103/PhysRevB.75.045315}.

\bibitem{Ado_2015}
I.~A. Ado, I.~A. Dmitriev, P.~M. Ostrovsky and M.~Titov,
\newblock \emph{\textit{Anomalous Hall effect with massive Dirac fermions}},
\newblock Europhys. Lett. \textbf{111}, 37004 (2015),
\newblock \doi{10.1209/0295-5075/111/37004}.

\bibitem{PhysRevLett.99.197202}
J.~Shi, G.~Vignale, D.~Xiao and Q.~Niu,
\newblock \emph{\textit{Quantum Theory of Orbital Magnetization and Its
  Generalization to Interacting Systems}},
\newblock Phys. Rev. Lett. \textbf{99}, 197202 (2007),
\newblock \doi{10.1103/PhysRevLett.99.197202}.

\bibitem{PhysRevB.86.214415}
G.~Zhu, S.~A. Yang, C.~Fang, W.~M. Liu and Y.~Yao,
\newblock \emph{\textit{Theory of orbital magnetization in disordered
  systems}},
\newblock Phys. Rev. B \textbf{86}, 214415 (2012),
\newblock \doi{10.1103/PhysRevB.86.214415}.

\bibitem{PhysRev.171.1065}
R.~M. Hornreich and S.~Shtrikman,
\newblock \emph{\textit{Theory of Gyrotropic Birefringence}},
\newblock Phys. Rev. \textbf{171}, 1065 (1968),
\newblock \doi{10.1103/PhysRev.171.1065}.

\bibitem{PhysRevB.82.245118}
A.~Malashevich and I.~Souza,
\newblock \emph{Band theory of spatial dispersion in magnetoelectrics},
\newblock Phys. Rev. B \textbf{82}, 245118 (2010),
\newblock \doi{10.1103/PhysRevB.82.245118}.

\bibitem{PhysRevLett.116.077201}
S.~Zhong, J.~E. Moore and I.~Souza,
\newblock \emph{Gyrotropic magnetic effect and the magnetic moment on the fermi
  surface},
\newblock Phys. Rev. Lett. \textbf{116}, 077201 (2016),
\newblock \doi{10.1103/PhysRevLett.116.077201}.

\bibitem{yoda2015current}
T.~Yoda, T.~Yokoyama and S.~Murakami,
\newblock \emph{\textit{Current-induced orbital and spin magnetizations in
  crystals with helical structure}},
\newblock Sci. Rep. \textbf{5}, 12024 (2015),
\newblock \doi{10.1038/srep12024}.

\bibitem{RevModPhys.87.1213}
J.~Sinova, S.~O. Valenzuela, J.~Wunderlich, C.~H. Back and T.~Jungwirth,
\newblock \emph{\textit{Spin Hall effects}},
\newblock Rev. Mod. Phys. \textbf{87}, 1213 (2015),
\newblock \doi{10.1103/RevModPhys.87.1213}.

\end{thebibliography}

\documentclass[submission, Phys]{SciPost}

\hypersetup{
    colorlinks,
    breaklinks=true,
    linkcolor={red!50!black},
    citecolor={blue!50!black},
    urlcolor={blue!80!black}
}

\usepackage{url}%comment this if you don't want doi links with breaks

\usepackage[bitstream-charter]{mathdesign}
\usepackage{times} 
\usepackage{graphicx} 
\usepackage{bm} 
\usepackage{amsmath} 
\usepackage{accents}
\usepackage{mathtools}
\usepackage{makecell}%for linebreaks in tables
\catcode`@=11

\def\nvphantom{\v@true\h@false\nph@nt}
\def\nhphantom{\v@false\h@true\nph@nt}
\def\nphantom{\v@true\h@true\nph@nt}
\def\nph@nt{\ifmmode\def\next{\mathpalette\nmathph@nt}%
  \else\let\next\nmakeph@nt\fi\next}
\def\nmakeph@nt#1{\setbox\z@\hbox{#1}\nfinph@nt}
\def\nmathph@nt#1#2{\setbox\z@\hbox{$\m@th#1{#2}$}\nfinph@nt}
\def\nfinph@nt{\setbox\tw@\null
  \ifv@ \ht\tw@\ht\z@ \dp\tw@\dp\z@\fi
  \ifh@ \wd\tw@-\wd\z@\fi \box\tw@}
\newcommand{\be}{\begin{equation}} 
\newcommand{\e}{\end{equation}} 
\newcommand{\bb}{\boldsymbol}
 
\DeclareMathOperator{\tr}{Tr}

\DeclareMathOperator{\sign}{sign}

\DeclareSymbolFont{usualmathcal}{OMS}{cmsy}{m}{n}
\DeclareSymbolFontAlphabet{\mathcal}{usualmathcal}

\begin{document}

\begin{center}{\Large \textbf{
Orbital Edelstein effect from the gradient of a scalar potential\\
}}\end{center}

\begin{center}
I.\,A.~Ado\textsuperscript{1$\star$},
M.\,Titov\textsuperscript{1},
Rembert A.\,Duine\textsuperscript{2,3} and
Arne Brataas\textsuperscript{4}
\end{center}

\begin{center}
{\bf 1} Radboud University, Institute for Molecules and Materials,\\Heyendaalseweg 135, 6525 AJ Nijmegen, The Netherlands
\\
{\bf 2} Institute for Theoretical Physics, Utrecht University,\\ Princetonplein 5, 3584 CC Utrecht, The Netherlands
\\
{\bf 3} Department of Applied Physics, Eindhoven University of Technology, \\P.O. Box 513, 5600 MB Eindhoven, The Netherlands
\\
{\bf 4} Center for Quantum Spintronics, Department of Physics, Norwegian University of Science and Technology, H\o gskoleringen 5, 7491 Trondheim, Norway
\\
${}^\star$ {\small \sf iv.a.ado@gmail.com}
\end{center}

\begin{center}
\today
\end{center}

% For convenience during refereeing (optional),
% you can turn on line numbers by uncommenting the next line:
%\linenumbers
% You should run LaTeX twice in order for the line numbers to appear.

\section*{Abstract}
{\bf
We study the orbital Edelstein effect (OEE) that originates from a particular inversion symmetry breaking mechanism: an asymmetric scalar potential. We compute OEE of this kind with the help of the Kubo formula in the diffusive regime for a parabolic band Hamiltonian. We also present a qualitative derivation of the effect. Both approaches give the same result. This result does not rely on spin-orbit coupling (SOC) and scales as a cube of the momentum relaxation time. In sufficiently clean large systems with weak SOC, OEE of this nature should exceed the spin Edelstein effect by orders of magnitude. It may also provide an alternative interpretation for some experiments concerning the spin Hall effect.
}

% Guideline: if your paper is longer that 6 pages, include a TOC
% To remove the TOC, simply cut the following block
\vspace{10pt}
\noindent\rule{\textwidth}{1pt}
\tableofcontents\thispagestyle{fancy}
\noindent\rule{\textwidth}{1pt}
\vspace{10pt}

%========================================================
%========================================================
\section{Introduction}
In nonmagnetic time-reversal invariant systems, an external electric field can generate nonequilibrium magnetization. If only the spin magnetic moments contribute to the latter, the corresponding effect is usually called the spin Edelstein effect (SEE)~\cite{EDELSTEIN1990233}. In the past decades, it attracted significant attention as a way to manipulate magnetic properties of materials by electrical means, in particular in the form of spin-orbit torques~\cite{RevModPhys.91.035004, 9427163}. The later terminology reflects the fact that SEE is a relativistic effect and requires spin-orbit coupling (SOC).

In recent years, the focus of research partly shifted from spin to orbital contributions to different magnetic phenomena, both equilibrium and nonequilibrium~\cite{PhysRevB.94.121114, PhysRevResearch.3.013275, atencia2024orbital}. Importantly, nonequilibrium magnetic effects of the orbital nature are often believed to be stronger than their spin counterparts as the former do not necessarily rely on SOC. The orbital Edelstein effect (OEE) represents a particular example of such an effect~\cite{doi:10.1021/acs.nanolett.7b04300, salemi2019orbitally}. It originates from the reaction of the orbital magnetic moments of the particles (rather than the spin ones) to the external electric field.

For finite OEE, inversion symmetry of the system should be broken~\cite{PhysRevResearch.3.013275, PhysRevB.107.115402}. In the past, OEE was studied for the following symmetry breaking mechanisms: chiral hoppings~\cite{doi:10.1021/acs.nanolett.7b04300}, \textit{sp}~hybridization at metal surfaces~\cite{go2017toward}, bilayer system asymmetry~\cite{PhysRevResearch.5.043294}, interaction-induced charge density wave order~\cite{PhysRevB.100.075136}, quadrupolar symmetry breaking~\cite{ishitobi2019magnetoelectric}, and, possibly, others. The more general magnetoelectric effect was also computed for asymmetric scattering from impurities~\cite{levitov1985magnetoelectric, PhysRevB.96.035120}.

However, one particular source of asymmetry that is very common in spintronics has been overlooked in the studies of OEE so far, namely the asymmetric scalar potential. To be more precise, possible effects of such a potential on OEE were considered only via the SOC term generated by it, while the potential itself was disregarded. At the same time, its gradient translates into an internal electric field that can be large both in bulk materials~\cite{ishizaka2011giant, PhysRevB.90.195210} and at interfaces~\cite{PhysRevB.55.16293, PhysRevLett.84.6074}. This internal field, as we demonstrate in the present work, gives rise to a strong OEE and cannot be ignored in this regard.

Below, we consider an infinite three-dimensional conducting system which is translationally invariant in the $xy$ plane and is modeled by a parabolic band Hamiltonian
\begin{equation}
\label{Ham}
    H=\frac{p^2}{2m}+U(z), \qquad p^2=p_x^2+p_y^2+p_z^2, \qquad p_\alpha=-i\hbar\nabla_\alpha,
\end{equation}
with an isotropic effective mass $m$. For simplicity, the internal asymmetric scalar potential $U(z)$ is approximated by a linear function\footnote{\label{footnote}In this formulation, $U(z)$ is unbounded at large $z$. Formally, this will make the particle density negative at some point. To avoid this, $U(z)$ can be understood as a bounded function approximated by $\delta'(q_z)$ in the Fourier space (here $\delta$ is the delta function). For example, we may use the standard recipe $U(z)\propto (\sin{zs})/s$ with sufficiently small $s$.}:
\begin{equation}
\label{V}
    U(z)=\frac{\partial U}{\partial z}z, \qquad \frac{\partial U}{\partial z}=const.
\end{equation}
A time-dependent external electric field $E_x$ is applied to this system along the $x$ direction and OEE is recognized as a linear in $E_x$ contribution $M_y$ to the nonequilibrium average density of the orbital magnetic moment operator in the $y$ direction.% We compute $M_y$ in the diffusive regime, thus the linear approximation for $U(z)$ is allowed.

The chosen model can be interpreted, e.g., as a simplified description of a bulk Rashba material~\cite{Rashba, PhysRevLett.5.371, ishizaka2011giant, PhysRevB.90.195210, PhysRevB.95.195164}, in which we completely ignore SOC. For our purposes, the most important characteristic of the system is the finite gradient of the scalar potential in Eqs.~(\ref{Ham}),~(\ref{V}), which any Rashba material provides. We also assume that $\partial U/\partial z\neq 0$ translates into a finite gradient of the particle density. The latter is just the condition of equilibrium with respect to the $z$ direction.

As we demonstrate below, the correct result for OEE in the considered model can be obtained using the following intuitive picture. Assume that the external field $E_x$ forces electrons to move with the drift velocity $v_x=v_{\text{drift}}$. The induced orbital magnetization $M_y$ is then proportional to the average value of $z v_x$. If this value is understood as a spatial average over a characteristic volume, weighed with the particle density $n(z)$, the gradient of the latter in the $z$ direction ensures that an uncompensated net orbital magnetization is created. The detailed version of this qualitative analysis is presented in Sec.~\ref{Elem_sec}. A full microscopic computation of OEE is performed in Sec.~\ref{Micro_sec}. Position independence of the computed orbital magnetization is established in Sec.~\ref{Spatial_sec}. In Sec.~\ref{Disco_sec}, we discuss the results.

%========================================================
%========================================================
\section{Elementary derivation}
\label{Elem_sec}
%========================================================
\subsection{Density distribution}
First, let us establish the relation between the density distribution and the scalar potential gradient. The equilibrium condition for the system is
\begin{equation}
    \mu(z)+U(z)=\mu(0)=\mu_0=const,
\end{equation}
where $\mu_0$ is the electrochemical potential. Note that the chemical potential $\mu(z)$ varies in the $z$~direction and that the Fermi-Dirac distribution function depends on $\mu_0$ rather than on $\mu(z)$:
\begin{equation}
\label{zeroT}
    f_\varepsilon=\Theta(\mu_0-\varepsilon),
\end{equation}
where we have assumed zero temperature. We introduce the retarded and advanced Green's functions for the Hamiltonian of Eq.~(\ref{Ham}), $G^{R,A}=\left(\varepsilon\pm i0-H\right)^{-1}$, and expand them with respect to the scalar potential gradient:
\begin{equation}
\label{G_exp}
    G^{R,A}=G^{R,A}_0+\frac{\partial U}{\partial z}G^{R,A}_0 \,z\, G^{R,A}_0, \qquad G^{R,A}_0=\left(\varepsilon\pm i0-H_0\right)^{-1}, \qquad H_0=\frac{p^2}{2m}.
\end{equation}
Here $H_0$ describes the homogeneous system and we disregarded higher-order terms in the Dyson series.

The particle density at the position $\bb r$ reads
\begin{equation}
\label{n_def}
    n(\bb r)=\frac{\gamma_s}{2\pi i}\int d\varepsilon f_\varepsilon\left[G^A(\bb r,\bb r)-G^R(\bb r,\bb r)\right],
\end{equation}
where $\gamma_s=2$ is the spin degeneracy factor. Next, we use the expansion of Eq.~(\ref{G_exp}) and the fact that $G^{R,A}_0$ correspond to a fully translationally invariant system. This allows us to switch to the momentum representation in Eq.~(\ref{n_def}),
\begin{equation}
    n(\bb r)=\frac{\gamma_s}{2\pi i}\int\frac{d^3p}{(2\pi\hbar)^3}d\varepsilon f_\varepsilon\left(\left\{G^A_0(\bb p)-G^R_0(\bb p)\right\}+z\frac{\partial U}{\partial z}\left\{\left[G^A_0(\bb p)\right]^2-\left[G^R_0(\bb p)\right]^2\right\}\right),
\end{equation}
with $G^{R,A}_0(\bb p)=1/\left(\varepsilon\pm i0-p^2/2m\right)$. Computing the momentum integrals by the residue theorem and employing Eq.~(\ref{zeroT}) to integrate over energy, we find for the particle density:
\begin{equation}
\label{n}
    n(\bb r)=n_0\left(1-\frac{3}{2}\frac{z}{\mu_0}\frac{\partial U}{\partial z}\right), \qquad n_0=n(0)=\frac{\gamma_s\sqrt{2(m\mu_0)^{3}}}{3\pi^2\hbar^3}.
\end{equation}
The expression for $n_0$ is of course well-known.

%========================================================
\subsection{Magnetization as a spatial average}
To compute the orbital magnetization, we need to define the orbital magnetic moment operator. The Hamiltonian of Eq.~(\ref{Ham}) in the presence of a magnetic field $\bb B=B_y \bb e_y$ $\mathcal B$ can be expressed as
\begin{equation}
    H_B=\frac{1}{2m}\left(\bb p-\frac{e}{c}\bb e_x B_y z\right)^2+U(z),
\end{equation}
where $e$ is the electron charge, $c$ is the speed of light, $\bb p=-i\hbar\bb\nabla$, and we disregarded the Zeeman term since it is not important for OEE. The $y$ component of the orbital magnetic moment operator then reads
\begin{equation}
\label{Mydef}
    \mathcal{M}_y=-\frac{\partial H_B}{\partial B_y}=-\frac{2\mu_B m_e}{\hbar} z v_x, \qquad \mu_B=-\frac{e\hbar}{2m_e c},
\end{equation}
where $\mu_B$ is the Bohr magneton, $v_x=p_x/m$, and $m_e$ is the electron mass in vacuum.

We would like to compute the average density of the observable $\mathcal{M}_y$. The result does not depend on the position (see Sec.~\ref{Spatial_sec}), and thus we compute it at $\bb r=0$. We assume that only the regions at distances of the order of the mean free path $l$ contribute to the result because at larger distances the coherence is lost due to scattering from impurities. For $v_x$, we substitute the drift velocity $v_{\text{drift}}=eE_x\hbar\tau/m$, where $\tau$ is the transport relaxation time measured in the inverse energy units. In this work, we consider isotropic scattering from delta-correlated scalar impurities, hence $\tau$ coincides with the scattering time. Under the assumptions made and using Eq.~(\ref{Mydef}), we~write 
\begin{equation}
\label{Mav}
    M_y=\frac{[\mathcal{M}_y]_{V_l}}{V_l}=-\frac{2\mu_B m_e}{\hbar}\frac{eE_x\hbar\tau}{m} \frac{1}{V_l}\int\limits_{V_l}d^3 r\, n(z)z ,%=-\frac{2\mu_B m_0}{\hbar}\frac{eE_x\hbar\tau}{m} \frac{1}{2l}\int\limits_{-l}^{l}d z\, z n(z),
\end{equation}
where $[\mathcal{M}_y]_{V_l}$ denotes the total orbital magnetic moment ``contained'' in the cube $V_l$. The latter is centered at $\bb r=0$, and the length of its edges is equal to $2l$. In Eq.~(\ref{Mav}), $d^3 r\,n(z)$ expresses the total number of particles in the volume $d^3 r$. Therefore, $M_y$ in this formula has a meaning of the orbital magnetization averaged over all particles in $V_l$.

Once Eq.~(\ref{n}) is substituted in Eq.~(\ref{Mav}), we integrate over $z$ and find
\begin{equation}
\label{Myelemfinal1}
    M_y=\frac{\mu_B m_e}{\hbar}\frac{eE_x\hbar\tau}{m}\frac{n_0}{\mu_0}\frac{\partial U}{\partial z}l^2.
\end{equation}
This result can also be expressed in terms of the charge current density $j_x=e n_0(eE_x\hbar\tau/m)$:
\begin{equation}
\label{Myelemfinal2}
    M_y=\frac{\mu_B m_e}{e\hbar}j_x\frac{l^2}{\mu_0}\frac{\partial U}{\partial z}=-\frac{1}{2c}j_x\frac{l^2}{\mu_0}\frac{\partial U}{\partial z}.
\end{equation}
We note that, since $l\propto\tau$ and $j_x\propto\tau$, the computed $M_y$ is of the third order with respect to $\tau$.

%========================================================
%========================================================
\section{Microscopic derivation}
\label{Micro_sec}
%========================================================
\subsection{Self-energy}
Let us now derive the results of Eqs.~(\ref{Myelemfinal1}),~(\ref{Myelemfinal2}) microscopically. For this, we supplement the Hamiltonian of Eq.~(\ref{Ham}) with a scalar Gaussian delta-correlated disorder potential with zero mean,
\begin{equation}
    \label{Ham_dis}
    H_{\text{dis}}=\frac{p^2}{2m}+U(z)+U_{\text{dis}}(\bb r), \quad \langle U_{\text{dis}}(\bb r_1)  U_{\text{dis}}(\bb r_2) \rangle=\alpha_{\text{dis}}\delta(\bb r_1-\bb r_2), \quad \langle U_{\text{dis}}(\bb r)\rangle=0,
\end{equation}
where the angular brackets denote the ensemble average. In the dc limit, at zero temperature, we express the orbital magnetization $M_y$ with the help of the Kubo-type~\cite{kubo1957statistical} formula,
\begin{equation}
\label{KUBOBUBO}
    M_y=\gamma_s\frac{e E_x\hbar}{2\pi V}\left\langle\tr{\left[v_x G^A_{\text{dis}} \mathcal{M}_y G^R_{\text{dis}}\right]}\right\rangle, \qquad G^{R,A}_{\text{dis}}=\left(\mu_0\pm i0-H_{\text{dis}}\right)^{-1},
\end{equation}
where $V$ is the system volume, $v_x=p_x/m$, and, due to $\partial f_\varepsilon/\partial\varepsilon=-\delta(\varepsilon-\mu_0)$, the energy argument of the Green's functions' is fixed at $\mu_0$ (from now on).

We compute $M_y$ in the noncrossing approximation~\cite{PhysRevB.75.045315, Ado_2015} as we are interested only in the contributions of the leading order with respect to $\tau$\footnote{For the same reason we neglected all contributions in Eq.~(\ref{KUBOBUBO}) that upon averaging become subleading for $\tau\to\infty$.}. In this case, averaging in Eq.~(\ref{KUBOBUBO}) boils down to adding diffusion ladder vertex corrections to either $v_x$ or $\mathcal{M}_y$ and to replacing each Green's function with its average. For our simple model, all vertex corrections vanish\footnote{Technically, they vanish as a result of the integration over the angles of momentum.}, thus we only need to average the Green's functions. In the first Born approximation, we have
\begin{equation}
    \langle G^{R,A}_{\text{dis}} \rangle=\left(\mu_0-\Sigma^{R,A}-H\right)^{-1}, \qquad \Sigma^{R,A}(\bb r)=\alpha_{\text{dis}}G^{R,A}(\bb r,\bb r),
\end{equation}
or, neglecting the real part of the self-energy $\Sigma^{R,A}$,
\begin{equation}
    \Sigma^{R}(\bb r)=-\Sigma^{A}(\bb r)=\frac{1}{2}\left[\Sigma^{R}(\bb r)-\Sigma^{A}(\bb r)\right]=\frac{\alpha_{\text{dis}}}{2}\left[G^{R}(\bb r,\bb r)-G^{A}(\bb r,\bb r)\right].
\end{equation}
Comparing this with Eq.~(\ref{n_def}) and using Eq.~(\ref{n}), we find
\begin{equation}
\label{n_diff}
    \Sigma^{R}(\bb r)=-\frac{i\pi\alpha_{\text{dis}}}{\gamma_s}\frac{\partial n(\bb r)}{\partial \mu_0}=-\frac{i}{2\tau}\left(1-\frac{z}{2\mu_0}\frac{\partial U}{\partial z}\right), \qquad \tau=\frac{\pi\hbar^3}{\alpha_{\text{dis}}\sqrt{2m^{3}\mu_0}}.
\end{equation}
In the following, we disregard the position-dependent part of the self-energy because its contribution to $M_y$ is smaller by the factor $1/(\mu_0\tau)$ than the leading order one. Hence the ensemble average of the Green's functions is provided by the expression
\begin{equation}
    g^{R,A}=\langle G^{R,A}_{\text{dis}} \rangle=\left(\mu_0\pm \frac{i}{2\tau}-H\right)^{-1},
\end{equation}
for which we introduced a short-handed notation $g^{R,A}$.

%========================================================
\subsection{Magnetization as a statistical average}
We can now see that, after averaging, the Kubo formula of Eq.~(\ref{KUBOBUBO}) reads
\begin{equation}
\label{KUBOBUBO2}
    M_y=\gamma_s\frac{e E_x\hbar}{2\pi V}\tr{\left[v_x g^A \mathcal{M}_y g^R\right]}.
\end{equation}
To evaluate it, one should use an expansion analogous to that of Eq.~(\ref{G_exp}),
\begin{equation}
\label{g_exp}
    g^{R,A}=g^{R,A}_0+\frac{\partial U}{\partial z}g^{R,A}_0 \,z\, g^{R,A}_0, \qquad g^{R,A}_0=\left(\mu_0\pm \frac{i}{2\tau}-H_0\right)^{-1}, \qquad H_0=\frac{p^2}{2m}.
\end{equation}
We substitute Eq.~(\ref{g_exp}) into Eq.~(\ref{KUBOBUBO2}) and employ the definition of Eq.~(\ref{Mydef}). The part independent of $\partial U/\partial z$ vanishes in the obtained expression as it corresponds to an inversion symmetric system. We are therefore left with
\begin{equation}
\label{KUBOBUBO3}
    M_y=-\gamma_s\frac{2\mu_B m_e}{\hbar}\frac{e E_x\hbar}{2\pi V}\frac{\partial U}{\partial z}\tr{\left[v_x g^A_0 z g^A_0 z v_x g^R_0\right]}+\text{c.c.} %+ v_xg^A_0 z v_x g^R_0 z g^R_0
\end{equation}

To evaluate Eq.~(\ref{KUBOBUBO3}), we regularize both position operators in it by means of the substitution $z\to(\sin{zs})/s$, with $s$ to be sent to zero in the end\footnote{See also footnote~\ref{footnote}.}. The same substitution is, in fact, used to derive the ``modern theory of orbital magnetization''~\cite{PhysRevLett.99.197202}. Although the usage is indirect in that case because the sine is introduced from the start to define the vector potential. As a result, position operators never appear in the derivations that follow. Technically, however, it leads to the same principal outcome: in momentum space, the position operator~$\bb r$ acts as a derivative with respect to~$\bb p$\footnote{Position operators under the trace in Eq.~(\ref{KUBOBUBO3}) must stand between two Green's functions. Moreover, in the momentum representation, one cannot use the cyclic permutation property to place a position operator in front of all other operators under the trace. Indeed, that would differentiate the entire integrand over the variable of integration and effectively nullify the result.}. For the ``modern theory'' of orbital magnetization, this can be seen in its Green's functions' formulation~\cite{PhysRevB.86.214415}.

Using the momentum space representation $i\hbar(\partial/\partial\bb p)$ of the position operator $\bb r$, we express the trace in Eq.~(\ref{KUBOBUBO3}) as an integral over momentum, integrate by parts twice, and arrive at
\begin{equation}
\label{KUBOBUBO4}
    M_y=\gamma_s\frac{2\mu_B m_e}{\hbar}\frac{e E_x\hbar}{2\pi}\frac{\partial U}{\partial z}\hbar^2\int \frac{d^3p}{(2\pi\hbar)^3}v_x^2(\bb p) g^R_0(\bb p) \frac{\partial}{\partial p_z}\left[g^A_0(\bb p)\frac{\partial g^A_0(\bb p)}{\partial p_z}\right]+\text{c.c.},
\end{equation}
where $g^{R,A}_0(\bb p)=1/\left(\mu_0\pm i/(2\tau)-p^2/2m\right)$. We also took into account that $v_x$ does not depend on $p_z$ and that all operators inside the integral commute with each other. With the help of partial integration and the relations $g^R_0g^A_0=i\tau(g^R_0-g^A_0)$ and $\partial g_0^{R,A}/\partial p_z=g_0^{R,A}v_z g_0^{R,A}$, one can bring Eq.~(\ref{KUBOBUBO4}) to the form
\begin{equation}
\label{KUBOBUBO5}
    M_y=\gamma_s\frac{2\mu_B m_e}{\hbar}\frac{e E_x\hbar}{2\pi}\frac{\partial U}{\partial z}\frac{\hbar^2\tau^2}{m}\int \frac{d^3p}{(2\pi\hbar)^3}v_x^2(\bb p) g^R_0(\bb p)g^A_0(\bb p),
\end{equation}
where several contributions of subleading orders with respect to $\tau$ have been omitted.

The integral here equals $2\pi n_0\tau/(m\gamma_s)$ which is, up to a prefactor, the longitudinal Drude conductivity. Therefore,
\begin{equation}
\label{final}
    M_y=\frac{\mu_B m_e}{\hbar}\frac{e E_x\hbar\tau}{m}\frac{\partial U}{\partial z}\frac{\hbar^2\tau^2}{m/2} n_0.
\end{equation}
Multiplying both the numerator and the denominator of the last fraction by the square of the Fermi velocity $v_F^2$, we reproduce the result of Eq.~(\ref{Myelemfinal1}) since $\mu_0=m v_F^2/2$ and $l=v_F\hbar\tau$.

%========================================================
%========================================================
\section{Position (in)dependence of the result}
\label{Spatial_sec}

%========================================================
\subsection{Expansions at an arbitrary $z=\widetilde{z}$}
In Eqs.~(\ref{G_exp}) and (\ref{g_exp}), we expanded the Green's functions $G^{R,A}$ and $g^{R,A}$ with respect to $(\partial U/\partial z)z$. Consequently, the Taylor expansions that we obtained for $n(\bb r)$ and $\Sigma(\bb r)$ were centered at $z=0$. The zeroth-order terms of these expansions, $n_0$ and $\tau$, explicitly enter the expression for $M_y$, Eq.~(\ref{final}). It is thus important to see what changes if the expansions for $n(\bb r)$ and $\Sigma(\bb r)$ are computed at an arbitrary $z=\widetilde{z}$.

Writing $U(z)=U(\widetilde{z})+(\partial U/\partial z)(z-\widetilde{z})$ and expanding the Green's functions in Eq.~(\ref{n_def}) with respect to $(\partial U/\partial z)(z-\widetilde{z})$, we find for the density distribution:
\begin{equation}
\label{n_tilde}
    n(\bb r)=\widetilde{n}\left(1-\frac{3}{2}\frac{z-\widetilde{z}}{\widetilde{\mu}}\frac{\partial U}{\partial z}\right), \qquad \widetilde{n}=n(\widetilde{z})=\frac{\gamma_s\sqrt{2(m\widetilde{\mu})^{3}}}{3\pi^2\hbar^3},
\end{equation}
where $\widetilde{\mu}=\mu_0-U(\widetilde{z})=\mu(\widetilde{z})$. For the self-energy, we can still use Eq.~(\ref{n_diff}), therefore
\begin{equation}
\label{n_diff_tilde}
    \Sigma^{R}(\bb r)=-\frac{i}{2\widetilde{\tau}}\left(1-\frac{z-\widetilde{z}}{2\widetilde{\mu}}\frac{\partial U}{\partial z}\right), \qquad \widetilde{\tau}=\frac{\pi\hbar^3}{\alpha_{\text{dis}}\sqrt{2m^{3}\widetilde{\mu}}}.
\end{equation}
Consequently, the Green's functions in Eq.~(\ref{KUBOBUBO2}) expand as
\begin{equation}
\label{g_exp_tilde}
    g^{R,A}=\widetilde{g}^{R,A}_0+\frac{\partial U}{\partial z}\widetilde{g}^{R,A}_0 \,(z-\widetilde z)\, \widetilde{g}^{R,A}_0, \qquad \widetilde{g}^{R,A}_0=\left(\widetilde\mu\pm \frac{i}{2\widetilde\tau}-\frac{p^2}{2m}\right)^{-1},
\end{equation}
which leads to the same formula for $M_y$ as that of Eq.~(\ref{KUBOBUBO4}), but with $g^{R,A}_0$ replaced by $\widetilde{g}^{R,A}_0$. Evaluating this formula, we find for the orbital magnetization:
\begin{equation}
\label{final_tilde}
    M_y=\frac{\mu_B m_e}{\hbar}\frac{e E_x\hbar\widetilde{\tau}}{m}\frac{\partial U}{\partial z}\frac{\hbar^2\widetilde\tau^2}{m/2} \widetilde{n}.
\end{equation}
Clearly, the obtained expression generalizes Eq.~(\ref{final}) and reduces to the latter if $\widetilde{z}=0$.

Since $\widetilde{\tau}=\tau(\widetilde{z})$ and $\widetilde{n}=n(\widetilde{z})$ both enter Eq.~(\ref{final_tilde}), it may seem that the result for $M_y$ depends either on~$\widetilde{z}$ or, which is worse, on the way we compute it. This, however, is not the case. The alleged dependence is determined by the product $\widetilde{\tau}^3\widetilde{n}$, which is actually $\widetilde{z}$-independent. Indeed, it can be seen from Eqs.~(\ref{n_tilde}),~(\ref{n_diff_tilde}) that $\widetilde{\mu}$ in $\widetilde{\tau}^3\widetilde{n}$ cancels out, while the remaining factors in this product are position-independent constants. Hence it does not matter at which point in space we expand the density and the self-energy: the result for $M_y$ is the same.

%========================================================
\subsection{Linear density of the orbital magnetic moment}
We derived the Kubo formula of Eq.~(\ref{KUBOBUBO4}) by considering the response of the operator $z v_x$ to the electric field $E_x$. This operator determines the full orbital magnetic moment of the system. Division of its response to $E_x$ by the total volume gives the nonequilibrium orbital magnetization~$M_y$ averaged over the entire system. It is not guaranteed however that the orbital moment is distributed uniformly in space. While the latter condition is obviously satisfied for the $xy$ plane, the distribution in the $z$ direction may, in principle, be nonuniform. Fortunately, the orbital magnetic moment at a particular $z=\widetilde{z}$ can be computed directly, with the help of another method that does not require any averaging over the $z$ direction.

Namely, instead of considering the full orbital magnetic moment of the system, we can consider its linear density. For this, we introduce the vector potential $\bb A=\bb e_x B_y[\sign{(z-\widetilde{z})}]/2$ that defines the magnetic field $\bb B=\bb e_y B_y \delta(z-\widetilde{z})$. The derivative of the Hamiltonian in the presence of such a potential with respect to $B_y$ (taken with a minus sign),
\begin{equation}
    \delta_z\mathcal{M}_y=-\frac{\partial}{\partial B_y}\left[\frac{1}{2m}\left(\bb p-\frac{e}{2c}\bb e_x B_y\sign{(z-\widetilde{z})}\right)^2\right]=-\frac{\mu_B m_e}{\hbar}v_x\sign{(z-\widetilde{z})},
\end{equation}
defines the operator of density of the $y$ component of the orbital magnetic moment over the $z$ direction at the $z=\widetilde{z}$ plane\footnote{Here, the notation $\delta_z$ is used to indicate that the operator $\delta_z\mathcal{M}_y$ has a meaning of density per length in the $z$ direction.}.

Using the Fourier transform of the Heaviside step function, it is straightforward to demonstrate that the response of this operator to $E_x$ divided by the area of the $z=\widetilde{z}$ plane is expressed by the same Eq.~(\ref{KUBOBUBO4}) that we derived for the response of the total orbital magnetic moment operator $\mathcal{M}_y$\footnote{With $g^{R,A}_0$ replaced by $\widetilde{g}^{R,A}_0$.}. As we have shown in this paper, evaluation of Eq.~(\ref{KUBOBUBO4}) leads to a $z$-independent result (for the considered system). Therefore, we can conclude that the spatial distribution of the orbital magnetic moment is indeed uniform and averaging of $\mathcal{M}_y$ over the volume can by replaced by averaging of $\delta_z\mathcal{M}_y$ over the plane. Both methods give the same result which should be interpreted as uniform nonequilibrium orbital magnetization.

%========================================================
%========================================================
\section{Discussion}
\label{Disco_sec}
%========================================================
\subsection{More on the elementary derivation}
The derivation presented in Sec.~\ref{Elem_sec} is based on an intuitive picture. In a way, it replaces averaging of the product $z v_x$ by independent averaging of $z$ and $v_x$. Strictly speaking, such a procedure is not justified. Nevertheless, for the considered system, it does provide the correct result for the orbital magnetization, as we have shown by performing a controllable microscopic computation. We anticipate this to be a consequence of the absence of spin-orbit coupling in the system. 

Moreover, the agreement of the results suggests a particular physical interpretation that we would like to mention. Conventionally, orbital magnetization is a density of orbital magnetic moment. In the diffusive regime, as it seems from our intuitive picture, this density is determined by a particular characteristic volume set by the length scale of the order of the mean free path.

%========================================================
\subsection{Other methods}
It may be possible to compute OEE in the considered model using the theory of spatial dispersion~\cite{levitov1985magnetoelectric, PhysRev.171.1065, PhysRevB.82.245118, PhysRevLett.116.077201}. At the same time, it does not seem plausible that this can be done with the help of the ``Boltzmann approximation''~\cite{yoda2015current, doi:10.1021/acs.nanolett.7b04300}. Indeed, the latter produces results that are linear with respect to $\tau$, whereas our computation gives the cubic dependence.

%========================================================
\subsection{Choice of the gauge}
In our computation, we chose the $\mathcal{M}_y\propto z v_x$ gauge for the orbital magnetic moment operator and dismissed the $\mathcal{M}_y\propto x v_z$ gauge. This choice is determined by the spatial symmetry of the system, which is translationally invariant in the $xy$ plane and inhomogeneous in the $z$ direction due to the scalar potential $U=U(z)$. The $\mathcal{M}_y\propto x v_z$ gauge does not match this symmetry because of its explicit dependence on the $x$ coordinate. In our approach, averaging of the operator $x v_z$ over the infinite $xy$ plane is mathematically ambiguous and not well justified. 

%========================================================
\subsection{Interfacial orbital Edelstein effect from the density gradient}
In Sec.~\ref{Elem_sec}, we presented a qualitative interpretation of OEE originating from the density gradient. In this picture, the nonequilibrium orbital magnetization at a certain point in the material is generated by electrons at distances of the order of the mean free path $l$ away from this point. At larger distances, coherence is lost due to scattering. This ``diffusive'' OEE can be realized in samples with large enough sizes that exceed $l$.

We argue, however, that the same physics may occur in smaller samples, in particular at the interfaces and in heterostructures with the structural inversion asymmetry. In this case, $l^2$ in our formula for OEE would be replaced by a different squared length scale. Of course, in such systems, scattering from the boundaries can also contribute to OEE.

We note that, in large clean systems, OEE originating from the density gradient may, in principle, be characterised by colossal length scales.

%========================================================
\subsection{Comparison with other effects}
Generation of nonequilibrium magnetization by electric fields and charge currents in nonmagnetic materials has been extensively studied with relation to the spin Edelstein effect and the spin Hall effect (SHE). In experiment, SEE and SHE are usually probed using the Faraday and Kerr effects or by measuring torques on magnetization~\cite{RevModPhys.87.1213, RevModPhys.91.035004}. While it is not yet clear whether OEE can produce considerable torques on magnetization, its signature in the Faraday and Kerr effects should be observable. Since OEE is not a relativistic effect, while both SEE and SHE are, we anticipate that magnetization induced due to OEE in Rashba materials can be stronger as compared with that related to SEE and SHE. In particular, this should be the case in sufficiently clean systems, in which the $\tau^3$ scaling of OEE would dominate the dependence on $\tau$ of SEE and SHE.

%========================================================
%========================================================
\section{Conclusion}
We studied the orbital Edelstein effect originating from the gradient of the scalar potential that translates into the particle density gradient. We presented a microscopic derivation of the effect and a qualitative one. Both provide the same result which does not rely on spin-orbit coupling and scales as a cube of the momentum relaxation time. We argue that the orbital Edelstein effect of this nature can be very strong in Rashba and other materials and should not be ignored in experiments on the spin Edelstein effect and the spin Hall effect.

%========================================================
%========================================================
\section*{Acknowledgements}
We are grateful to Sergii Grytsiuk and Yuriy Mokrousov for informative discussions.

\paragraph{Funding information}
 R.\ A.\ D. has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant No.~725509). The Research Council of Norway (RCN) supported A.\ B. through its Centres of Excellence funding scheme, project number 262633, ``QuSpin'', and RCN project number 323766. M.\ T. has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sk\l{}odowska-Curie grant agreement No 873028.

%========================================================
%========================================================

%\begin{appendix}
%\section{}
%\end{appendix}
%========================================================
%========================================================
\bibliography{Bib.bib}

\nolinenumbers

\end{document}